OpenStudy (anonymous):

Last fall, a gardener planted 63 iris bulbs. She found that only 57 of the bulbs bloomed in the spring. •Find the empirical probability that an iris bulb of this type will bloom. Give answer as a fraction in lowest terms. •How many of the bulbs should she plant next fall if she would like at least 60 to bloom?

OpenStudy (dumbcow):

P = 57/63 57/63 = 60/x x = (63*60)/57 x = 66.3 she should plant 67

OpenStudy (dumbcow):

oops it reduces P=19/21

OpenStudy (anonymous):

what method did u use to get P=19/21

OpenStudy (dumbcow):

divide numerator and denominator by 3

OpenStudy (anonymous):

oh ok i will try that thanks.

OpenStudy (anonymous):

i got 19/22.1